A fixed mass of a gas occupies 100cm cubed ,at-15cm cubed and 650mmHg pressure at what temperature will it have a volume of 150cm cubed if the pressure is adjusted to 680 mmHg?

1
Expert's answer
2022-05-23T17:01:09-0400

V0 – initial volume = 100cm3 = 10dm3 = 10 l.

T0 – initial temperature = -15°C = 258 K.

P0 – initial pressure = 650 mm Hg.

V – final volume = 150cm3 = 15dm3 = 15 l.

P – final pressure = 680 mm Hg.

T – final temperature = ?


We derive the formula for finding T using the Mendeleev - Clapeyron equation:

PV = nRT –> n= PV/RT

n0 = P0 V0 / RT0

Since n = n0,

PV/RT = P0V0/RT0 –> T= PVT0/P0V0


T = PVT0 /P0V0 = 680 * 15 * 258 / 650 * 10 =

= 2631600/6500= 404,86 K

404,86 K => 131,86°C

ANSWER: 131,86°C ≈ 132°C